OpenStudy (anonymous):

Okay how would you factor x^2-49y^2

OpenStudy (kirbykirby):

(x+7)(x-7)

OpenStudy (anonymous):

thank you

OpenStudy (kirbykirby):

oops. (x-7y)(x+7y) I didn't see the y^2

OpenStudy (anonymous):

I figured that but thanks

OpenStudy (kirbykirby):

There's a property that says a^2-b^2 = (a+b)(a-b) here, a = x, b = 7y (square root of 49y^2 is 7y)

OpenStudy (anonymous):

@kirbykirby ho would you use that property on x^4-81y^4

OpenStudy (kirbykirby):

say a=x^2 and b = 9y^2 so, (x^2-9y^2)(x^2+9y^2)... and the first factor can be simplified more, = [(x+3)(x-3)](x^2+9^y2)

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

do i then factor out the x^2+9^y^2

OpenStudy (anonymous):

i meant x^2+9y^2

OpenStudy (kirbykirby):

Oh i forgot the y again, it's (x+3y)(x-3y)... sorry about that >_> But no you can't simplify x^2+9y^2. It's actually an irreducible polynomial so it can't be factored anymore.

OpenStudy (kirbykirby):

there is no factorization for the form (a^2+b^2)

OpenStudy (anonymous):

okay thank you again

OpenStudy (kirbykirby):

np